20 ml of a sample water on titration with N/50 H2SO4 required 9.6 ml of the acid to Phenolphthalein end point and 13 ml of the acid to Methyl orange end point. Determine the type and amount of alkalinity present in the water sample


1
Expert's answer
2021-10-14T04:04:24-0400

Here diacidic base is present which give two end point one with phenalphthein and second with methy orange


For phenolphthalein end point-


"\frac{1}{2}\ Meq\ of\ Diacidic\ base = Meq.\ of\ of \ 9.6ml \ \frac{N}{50}\ HCl"

="\frac{1\times 9.6}{50\times1000}"


="Meq.\ of\ Diacidic\ base=2\times Meq.\ of\ H2SO4"


For methyl orange end point-

"Meq.\ of\ monoacidic\ base\ formed=Meq.\ of\ 1ml of\ extra\ HCl\ used"


"=\frac{1\times1}{50\times 1000}"

= "\frac{1}{50000 }Meq."




Normality of diacidic alkalinity present in 20 ml water = "\frac{2\times 1000}{6225\times 20}=\frac{20}{6225}=\frac{4}{1245}\ N"